# Knowledge Distillation-Based Model Extraction Attack using Private Counterfactual Explanations

### v1

@@ 3.1 Counterfactual Explanations (CF)
$V_{\text{CounterGAN-bb}}(D,G)=\frac{\sum_{i}C_{t}(x_{i})\log D(x_{i})}{\sum_{i}C_{t}(x_{i})}+\frac{1}{N}\sum_{i}\log(1-D(x_{i}+G(x_{i})))+\text{Reg}(G,\{x_{i}\}),$ (1)
@@ 3.3 Differential Privacy
One widely utilized approach for addressing numerical inquiries involves the incorporation of randomized approaches that add calibrated random noise, which works by introducing sufficient noise to the input or output of the mechanism to obscure the potential contribution of any individual record in the data while simultaneously maintaining the overall accuracy of the analysis. In Def. 3.1, we detail the DP inequality.
Definition 3.1 (Differential Privacy)
@@
Def. 3.1 states that an algorithm $M$ ($\varepsilon,\delta$) is differentially private if, for all subsets $S$ of the range of $M$ and all pairs of inputs $x$ and $y$ differing by at most one data instance, the probability of $M(x)$ outputting $S$ is bounded by a factor of $e^{\varepsilon}$ and an error margin $\delta$.
@@
4 Scenario and Problem Formulation
5 Methodology
In this section, we first describe our proposed KD-based MEA and train a threat model $t_{\Upsilon}$. Then, we describe our proposed methods for integrating DP within the CF generation process.
5.1 Knowledge Distillation-based Model Extraction Attack
As an attacker, specifically, we query the model with random queries and collect the query output of the target model, along with the provided CFs (Step 3 in Fig. 2). After successfully collecting a dataset of CFs (Step 4 in Fig. 2), the attacker employs our KD-vases MEA approach on the collected dataset as input data to train $t_{\Upsilon}$. The rationale for selecting KD as the extraction method lies in its unique capability to address the problem not only as a standard supervised classification task but also in its proficiency in precisely imitating the probability distribution. Specifically, we propose to train $t_{\Upsilon}$ by minimizing the loss constituted by the threat model classification loss in addition to the distillation loss. To emphasize the importance of mimicking the output probabilistic distribution from the target model to the threat model, we set the distillation loss to the Jensen-Shannon (JS) 4 metric. The choice of JS is to tackle issues related to divergence and ensure a symmetric output, so we suppose that the attacker replaces the KL divergence metric with JS. Afterward, the attacker trains the threat model by minimizing the loss of KD (Step 5 in Fig. 2) until the convergence of the agreement on a separate validation set.
5.2 Private CFs Generation with DP
We are currently aligned with the service provider and are presenting a strategy to mitigate MEA by avoiding providing Private CFs. Our goal is to examine the effects of incorporating DP into the CF generation process on MEA performance and the quality of the provided CFs considering metrics such as actionability, predction gain, and realism, which as explained later in sec 6.4. Additionally, we aim to demonstrate how private CFs align with the distribution of the training set. We propose an approach to integrating DP into the CF generation process and generating private CFs. Note that our methodology does not memorize or expose statistical information about the training set. We propose to incorporate DP into the generator of the CounteRGAN. The primary objective is to prevent the generation of CFs that closely resemble the private training data, thereby reducing the resemblance of CFs with the training data.
6 Experimental Settings
In this section, we will discuss the baseline scenarios we considered, in addition to the configuration we followed to train our models and our experimental datasets.
6.1 Baseline Scenarios
We compare our proposed approach for performing MEA, KD-based MEA, to a baseline scenario, presented in [12], and referred to as Direct Train, which trains a threat model as a standard supervised classification task. We examine three scenarios pertaining to the MLaaS provider’s inclusion or exclusion of CF explanations alongside the predictions of the ML model:
  • No CF: In this scenario, the MLaaS does not provide the user with any CFs.
  • CF: In this case, the MLaaS provides CFs suing counterGAN.
  • Private CF: The MLaaS offers private CFs by implementing our proposed approach for generating private CFs, discussed in Sec. 5.
Applying each MEA approach to the three scenarios results in the following six scenarios:
  • KD-based Private CF. We employ our proposed KD-based approach on CFs extracted from the differentially private generator, and the focus shifts to examining the potential effect of integrating DP in the CF generation.
  • KD-based CF. We employ our proposed KD-based approach, trained with CFs, to understand the influence of KD when applied alongside CF.
  • KD-based No CF. We employ our proposed KD-based approach on the randomly generated data points, and the focus shifts to examining the potential enhancement when KD is utilized independently of CFs and analyzing the power of CFs alongside.
  • Direct No CF. The MEA is performed following the baseline [12]. We train a model directly on the randomly generated data points, and the focus shifts to examining the potential of direct queries and analyzing the power of KD and CFs.
  • Direct CF. We train a model directly on CFs, this dimension is dedicated to evaluating the standalone power of CFs without the incorporation of KD, providing insights into the capabilities of CFs separately.
  • Direct Private CF. We train a model directly on CFs extracted from the differentially private generator, the focus shifts to examining the potential of direct queries and to analyzing the power of KD and CFs alongside.
6.2 Models Training
6.2.1 Target Model
We adhere to consistent training procedures for the target model (to be deployed and queried by the API). We train a DNN comprising of 16 hidden layers to create a complex model. The layer configurations consist of 64, 32, 16, 32, 64, 128, 64, 32, 128, 64, 128, 64, 128, 64, 32, and 16 neurons in each layer, respectively. The activation function is set to gelu for all layers, while the output layer employs softmax to produce model outputs as confidence score probabilities. We specify the optimizer as the Adam optimizer to minimize the cross-entropy loss function, with default parameter initialization provided by Keras. Each target model is trained using 80% of the training dataset.
6.2.2 Threat model
6.2.3 CF generator
6.3 Datasets
@@
6.4 Evaluation metrics
Effectiveness of MEA: A commonly employed metric for assessing the effectiveness of MEA is the Agreement measure. Agreement measures the degree of alignment between two models, i.e., the similarity in the predictions between the two different ML models. In the context of MEA methods, agreements assess the output similarity between the predictions of an extracted model $t_{\Upsilon}$ and those of the target model $f_{\theta}$ for a given set of data records (see Eqn. 5). Since our goal is to replicate the behavior of a target model with an extracted model, a higher agreement means a more successful MEA, and hence, a more effective strategy for MEAs.
The Realism is a metric used to quantify how closely a data instance fits a known data distribution. Since in our work, we employ DP to avoid providing CF that memorizes or exposes statistical information about the training set, we employ realism to quantify how well CFs and private CFs fit in with the known data distribution. Inspired by the strategies proposed in [33, 22], we train a denoising autoencoder on the noised training set, and compute the realism of a data point as the reconstruction error as the mean squared error (MSE) of the autoencoder (Eq. 6). A lower value of $realism$ corresponds to a better fit of the data point within the data distribution. We employ this metric to assess the authenticity of CFs and to examine the impact of the data perturbation with DP on CF. Furthermore, realism is employed to compare a data point with its associated CF.
Actionability is a commonly used metric that allows to measure the quality of a CF. Actionability assumes both the number of altered features and the degree of those changes (amount of modification in a feature) in a CF relative to the original data instance. We compute Actionability by taking the L1 norm of the absolute difference between the data point and its corresponding CF (Eq. 7. A low value of actionability (desired) indicates that a relatively small subset of the input features has been perturbed, with respect to original data instance, to achieve a different outcome by the classifier, and hence, a low actionability suggests better quality CFs.
By computing this comprehensive set of metrics, our goal is to explore how DP impacts the feasibility of taking actionable steps and how this influence contributes to the prediction gain outcomes.
7 Results and Evaluation
7.1 Agreement for Model Extraction Attack Results
(a) GMSC Dataset
50100200300500100060606060707070708080808090909090Queries NumberAgreementKD-based No CFDirect No CF
(b) Credit Card Fraud Dataset
501002003005001000606060607070707080808080Queries NumberAgreementKD-based Private CFDirect Private CF
(c) Housing Dataset
Figure 3: The agreement values achieved by the MEA approach in the various scenarios across the a) GMSC dataset, b) Bank Fraud Dataset, c) Housing Dataset with respect to number of queries made.
Figure 3 show the agreement achieved by KD-based MEA and Direct in the various scenarios with respect to the number of queries made to the MLaaS API across the GMSC, the Credit Card Fraud, and the California Housing datasets. Results show a consistent pattern across all scenarios as agreement initially increases significantly as number of queries increase until a point where additional queries cease to yield a notable increase in agreement. This suggests that there is a saturation point in the MEA, and further queries may not necessarily contribute substantially to enhancing agreement of the attack.
We now focus on analyzing the effectiveness of exploiting CFs for performing MEAs, comparing the agreement achieved by each of KD-based MEA and Direct when CF explanations are used to that when only data points are used (i.e., comparing Direct No CF to Direct-CF, and KD-Based No CF to KD-based CF). In both cases, when CFs are exploited, the MEA is more effective, achieving an agreement significantly higher than its counterpart. For instance, in GMSC dataset, Direct-CF achieves an agreement ranging between 80 and 85 while Direct No CF achieves an agreement ranging between 55 and 75 (showing an additional agreement of up to 25%). For Credit Card Fraud dataset, the difference is significantly larger as Direct-CF shows an agreement ranging between 83 and 90 while Direct-no-CF has an agreement ranging between 58 and 75. This difference is also observed for California Housing dataset where exploiting CFs allows for an additional agreement of around 10%. Similarly, analyzing the agreement achieved when employing our proposed approach with CFs to its counterpart (i.e., KD-based CF vs. KD-Based No CF) across the three data sets, we notice that exploiting CFs permits obtaining up to additional 27% of agreement (for GMSC dataset at 50 queries) and, in the worst case, an additional 4% on agreement (for Housing dataset at 1000 queries). These results confirm previous findings in literature (e.g., [12] and [52]) which show the negative impact of using model’s CF explanations for performing MEA.
In the case when CFs are provided by the MLaaS and are then exploited by the attacker, i.e., KD-based CF vs. Direct CF, results show that across all the 3 datasets, the agreement achieve by KD-based CF agreement always surpasses the Direct-CF, independent of the number of queries. In the case KD-based CF maintains an additional 2% to 6% agreement across all datasets. Similarly, in the case where the MLaaS provide private CFs, i.e., comparing KD-based Private CF to Direct Private CF, KD-based Private CF maintains a slightly additional range from 1% to 2% in most query numbers across all datasets. The observed decrease in agreement of MEA with private CFs may be attributed to the privacy-preserving nature of DP. Differential privacy introduces intentional noise into the data to protect individual privacy, making it more challenging for attackers to accurately extract sensitive information. In the context of KD-based MEA, when DP is employed, the noise introduced during the training process to achieve privacy guarantees could interfere with the teacher-student model’s ability to transfer knowledge much more effectively.
In this consideration, using CFs significantly reduces the number of instances required for MEA to reach a specific high agreement level. For example, on the GMSC dataset, achieving 80% agreement with CFs took just 50 queries, while accomplishing the same level without CFs required over 1000 queries. This pattern holds across datasets, such as the Credit Card Fraud Dataset, where 83% agreement needed only 50 CF queries compared to over 1000 queries without CFs. In essence, the use of CFs consistently leads to a substantial reduction in the number of queries needed to achieve high agreement levels across diverse datasets. In addition to that, our proposed methodology KD-based MEA consistently results in higher agreement levels of MEA compared to the baseline approach. This holds in various scenarios and is independent of whether the MLaaS provides users with CF explanations, and across diverse datasets and query numbers. This analysis emphasizes the effectiveness of our proposed method, utilizing KD to achieve superior agreement levels. Importantly, when CFs are employed, MEA requires significantly fewer instances to reach a specific high agreement level.
This highlights the efficiency and efficacy of incorporating CFs in conjunction with KD to enhance the performance of the MEA.
We now analyze the impact of generating private CFs through incorporating DP in the explainer, i.e., by incorporating DP during the generation process of explanations, on mitigating MEA. Specifically, we aim to quantify the extent of protection offered by providing users with private CFs (as opposed to non-private CFs) such as to prevent attackers from leveraging vulnerabilities associated with non-private CFs. To this end, we compare the performance of each of KD-based CF and Direct CF to its private CF counterparts (KD-based Private CF and Direct Private CF, respectively). For the GMSC dataset, results show that for the relatively low number of queries (50 to 200), the difference is vast between each scenario with private CF and its counterpart. For instance, KD-based Private CF achieves an agreement ranging between 60 and 75 while that KD-based CF obtains an agreement between 83 and 87.
Moreover, for the GMSC dataset, we observe that the incorporation of DP shows that our proposed strategy can maintain agreement levels comparable to scenarios without CFs and less than when incorporating traditional CFs, with an agreement range of 60% to 83% with KD-based CF and 60% to 81% with direct training, compared to 57% to 79% with KD-based MEA no CF and a range of 57% to 76% with Direct No CF. For the credit card fraud dataset, a similar trend is observed. With KD-based Private CF, we achieve a range of 63% to 82% and 52% to 79% for a direct-DP train, compared to KD-based no CF of 60% to 81% and a range of 58 to 75 with Direct No CF. For the California housing dataset, the agreement of KD-based CF ranges from 59% to 68% with KD and 57% to 67% with Direct Private CF, compared to 59% to 75% with KD-based no CF, and direct no CF ranges from 57% to 74%. It is also worth noting that, KD-based Private CFis slightly better than direct training over all datasets.
Results suggest that the incorporation of DP can play a crucial role in maintaining agreement levels comparable to scenarios without CFs.
7.2 Impact of Incorporating DP in CF generator on quality of explanations
@@
Table 1 reports the prediction gain of CounteRGAN and the actionability of the CFs achieved with and without DP across the three datasets. Note that the focus lies on comparing the metrics on the CFs exclusively between scenarios with and without DP, regardless of the applied MEA approach. Results show that CounteRGAN achieves a higher prediction gain (a higher probability shift, desired) when DP is not incorporated within the explainer than when employed for GMSC (0.243 $\pm$ 0.011 versus 0.121 $\pm$ 0.01) and Bank Fraud datasets (0.700 $\pm$ 0.084 versus 0.445 $\pm$ 0.06) while for Housing dataset, CounteRGAN shows a prediction gain slightly higher in the case where DP is incorporated rather than when not (0.678 versus 0.633). These findings highlight that the integration of DP has an impact on prediction gain, constraining the explainer’s progress toward the desired class.
In terms of actionability which is measured as the degree of perturbation or modification of the CF compared to the initial point, the results show that the explainer generates CFs with lower values of actionability when DP is incorporated than when not across all datasets. For instance, for GMSC dataset, the actionability of CFs with DP is 16.981 ($\pm$ 0.158) while that without DP is 24.5 ($\pm$ 0.364). Similarly, for bank fraud and housing datasets, the actionability is reduced when incorporating DP from 35.269 $\pm$ 0.328 to 10.507 $\pm$ 0.238, and from 3.852 $\pm$ 0.053 to 1.004 $\pm$ 0.016 respectively. This means that that the practical usefulness of the CFs or perturbations in guiding actionable decisions is reduced, suggesting that the privacy constraints imposed might compromise the effectiveness of CF analyses in providing actionable guidance for influencing desired outcomes.
@@
Table 2 reports the realism of the data points used as initial queries and the CFs generated by the explainer across the three data sets. Results show that the Realism of CFs without incorporating DP in the generator is the lowest in comparison to the Realism of random points and that of CFs when DP is employed, meaning that the CF generator has produced more realistic data points (lower Realism), attempting to align the CFs of random points with the distribution of the training data. In particular, for the GMSC dataset, the randomly generated points have an average realism of 15.6 ($\pm$ 0.03), and the CF generator has produced corresponding CFs with a realism of 8.5 ($\pm$ 0.14). Similarly, for the Bank Fraud data set, the Realism of random points is 3.104 ($\pm$ 0.019) while for the corresponding CFs, it is slightly lower at 3.07 ($\pm$ 0.16), respectively. For the Housing data set, the realism of data points is 2.07 ($\pm$ 0.04) while their corresponding CFs had a realism of 1.35 ($\pm$ 0.03). This implies that the explainer generates CFs with lower realism from the original queried random points and aligns them more closely with the distribution of the training data. The private CFs preserve a realism very similar to that of random points. More specifically, for the GMSC dataset, the private CFs have preserved their initial average level of realism of 15.723 ($\pm$ 0.033). Similarly, the housing dataset preserves a realism of 2.0 ($\pm$ 0.01). For the bank fraud dataset, the realism of the CFs deviates by one degree from the distribution, resulting in a realism of 4.73 ($\pm$ 4.73). With the incorporation of DP, realism can be preserved indicating the effectiveness of the Private CFs in maintaining the quality and distribution of the original data. The findings might vary depending on the dataset, as shown by the slight deviation in realism for the Bank Fraud dataset.
@@
8 Conclusion

### v2

@@ A.1.1 Counterfactual Explanations (CF)
$\begin{split}V_{\text{CounterGAN-bb}}(D,G)=&\frac{\sum_{i}C_{t}(x_{i})\log D(x_{i})}{\sum_{i}C_{t}(x_{i})}\\ &+\frac{1}{N}\sum_{i}\log(1-D(x_{i}+G(x_{i})))\\ &+\text{Reg}(G,\{x_{i}\}),\end{split}$ (5)
@@ A.1.3 Differential Privacy
One widely utilized approach for addressing numerical inquiries involves the incorporation of randomized approaches that add calibrated random noise, which works by introducing sufficient noise to the input or output of the mechanism to obscure the potential contribution of any individual record in the data while simultaneously maintaining the overall accuracy of the analysis. In Def. A.1, we detail the DP inequality.
Definition A.1 (Differential Privacy)
@@
Def. A.1 states that an algorithm $M$ ($\varepsilon,\delta$) is differentially private if, for all subsets $S$ of the range of $M$ and all pairs of inputs $x$ and $y$ differing by at most one data instance, the probability of $M(x)$ outputting $S$ is bounded by a factor of $e^{\varepsilon}$ and an error margin $\delta$.
@@
A.2 Datasets
@@
A.3 Impact of Incorporating DP in CF generator on quality of explanations
@@
Table 1 reports the prediction gain of CounteRGAN and the actionability of the CFs achieved with and without DP across the three datasets. Note that the focus lies on comparing the metrics on the CFs exclusively between scenarios with and without DP, regardless of the applied MEA approach. Results show that CounteRGAN achieves a higher prediction gain (a higher probability shift, desired) when DP is not incorporated within the explainer than when employed for GMSC (0.243 $\pm$ 0.011 versus 0.121 $\pm$ 0.01) and Credit Fraud datasets (0.700 $\pm$ 0.084 versus 0.445 $\pm$ 0.06) while for Housing dataset, CounteRGAN shows a prediction gain slightly higher in the case where DP is incorporated rather than when not (0.678 versus 0.633). These findings highlight that the integration of DP has an impact on prediction gain, constraining the explainer’s progress toward the desired class.
In terms of actionability which is measured as the degree of perturbation or modification of the CF compared to the initial point, the results show that the explainer generates CFs with lower values of actionability when DP is incorporated than when not across all datasets. For instance, for GMSC dataset, the actionability of CFs with DP is 16.981 ($\pm$ 0.158) while that without DP is 24.5 ($\pm$ 0.364). Similarly, for credit fraud and housing datasets, the actionability is reduced when incorporating DP from 35.269 $\pm$ 0.328 to 10.507 $\pm$ 0.238, and from 3.852 $\pm$ 0.053 to 1.004 $\pm$ 0.016 respectively. This means that that the practical usefulness of the CFs or perturbations in guiding actionable decisions is reduced, suggesting that the privacy constraints imposed might compromise the effectiveness of CF analyses in providing actionable guidance for influencing desired outcomes.
@@
Table 2 reports the realism of the data points used as initial queries and the CFs generated by the explainer across the three data sets. Results show that the Realism of CFs without incorporating DP in the generator is the lowest in comparison to the Realism of random points and that of CFs when DP is employed, meaning that the CF generator has produced more realistic data points (lower Realism), attempting to align the CFs of random points with the distribution of the training data. In particular, for the GMSC dataset, the randomly generated points have an average realism of 15.6 ($\pm$ 0.03), and the CF generator has produced corresponding CFs with a realism of 8.5 ($\pm$ 0.14). Similarly, for the Credit Fraud data set, the Realism of random points is 3.104 ($\pm$ 0.019) while for the corresponding CFs, it is slightly lower at 3.07 ($\pm$ 0.16), respectively. For the Housing data set, the realism of data points is 2.07 ($\pm$ 0.04) while their corresponding CFs had a realism of 1.35 ($\pm$ 0.03). This implies that the explainer generates CFs with lower realism from the original queried random points and aligns them more closely with the distribution of the training data. The private CFs preserve a realism very similar to that of random points. More specifically, for the GMSC dataset, the private CFs have preserved their initial average level of realism of 15.723 ($\pm$ 0.033). Similarly, the housing dataset preserves a realism of 2.0 ($\pm$ 0.01). For the Credit fraud dataset, the realism of the CFs deviates by one degree from the distribution, resulting in a realism of 4.73 ($\pm$ 4.73). With the incorporation of DP, realism can be preserved indicating the effectiveness of the Private CFs in maintaining the quality and distribution of the original data. The findings might vary depending on the dataset, as shown by the slight deviation in realism for the Credit Fraud dataset.
@@
A.4 Limitations and Future Work
The limitations of our current approach and potential avenues for future research are the following:
  • Attack Scope: Our work focuses on MEA leveraging CFs, it is important to acknowledge the potential existence of other privacy attacks such as membership inference and model inversion that warrant further investigation. To this end, in future work, we plan to evaluate the effectiveness of our proposed private GAN-based CF generation approach against other attack types.
  • Privacy-Performance Trade-off: In this work, we made the first attempt to integrate DP in GAN-based CF generators. We assume that GAN-based CFs should be privacy-preserving against random and zero-knowledge queries. This means it should not reveal any realistic CFs when queried by random noise. A future direction is to thoroughly evaluate the impact of various DP levels on considering both their influence on mitigating various privacy attacks including MEA and their impact on the effectiveness and quality of CFs with DP.
  • Focus on Deep Learning Applications: KD is effective with DNNs due to its complex architecture and ability to learn rich data representations. Future research aims to explore adapting KD for other ML algorithms such as tree and ensemble-based algorithms and to potentially unlock similar performance enhancements.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We summarized our contributions in the abstract and introduction.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: The Limitations are discussed as a separate section in the appendix.
  9. 3.
    Theory Assumptions and Proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: All assumptions stated are referenced.
  13. 4.
    Experimental Result Reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: All the steps required for reproducibility are provided in the experimental settings.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: We use publicly available datasets (references to datasets are reported in the paper). We provide sufficient instructions and code to reproduce the main experimental results (libraries, parameters, etc.).
  21. 6.
    Experimental Setting/Details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: We provide all the necessary informations.
  25. 7.
    Experiment Statistical Significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We report the results of the average of 10 runs with randomly chosen subsets for each experiment of the MEA KD-based attack.
  29. 8.
    Experiments Compute Resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: Experiments were computed with a machine of intel Core i7, a GPU of RTX 3070, and 8 GB of RAM.
  33. 9.
    Code Of Ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We follow the ethics guideline.
  37. 10.
    Broader Impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [N/A]
  40. Justification: We discuss the effect of performing a model extraction attack, and we propose a mitigation method.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: The paper poses no such risks.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: We use publicly available datasets and codes from Kaggle that we referenced.
  49. 13.
    New Assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [N/A]
  52. Justification: We do not provide new assets.
  53. 14.
    Crowdsourcing and Research with Human Subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. Justification: We do not involve crowdsourcing nor research with human subjects.
  57. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  58. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  59. Answer: [N/A]
  60. Justification: The paper does not involve crowdsourcing nor research with human subjects.